OpenStudy (anonymous):

find a polynomial of degree 3 that has zeros 1,-1 and 4

OpenStudy (anonymous):

(x-1)(x+1)(x-4)

OpenStudy (anonymous):

\[(x-1)(x+1)(x-4)\] will do it. you have to multiply out probably

OpenStudy (anonymous):

X^3-4x^2-x+4

OpenStudy (anonymous):

how di du find that?

OpenStudy (anonymous):

Will, we are looking for zeros at -1,1, and 4 So, poylinomial is fairly easy to do, since we simply have to factor it and set each term equal to zero. A polynomial wiht zeros at -1,1, and 4 is given by (x+1)=0, (x-1)=0, and (x-4)=0

OpenStudy (anonymous):

in each case solve for x

OpenStudy (anonymous):

clear yes? factor theorem says that if r is a zero of p(x) then p(x) = (x-r)q(x) just factor all the way down